OpenStudy (anonymous):

Use a lemma to prove that if 5x+5 is even, 9x+3 is even. @Mertsj @myininaya @mathstudent55 @abb0t @SolomonZelman @wio @kelliegirl33 @Compassionate @Loser66 @eliassaab @Ashleyisakitty @malevolence19 @helpme1.2 @tester97 @tHe_FiZiCx99 @UsukiDoll @NoelGreco @Mr.ClayLordMath @Kristen17 @shelbygt520 @sourwing @tanya123 @adrynicoleb @thtshortchick @Savannah_Noelle @youngamerican @dariyana @kawaiicat123 @Hayden14 @brendalopez @UH60blackhawk @Kittta @Figureskater120 @canadalover_ @linh412986 @ComeAlongP0nd @Taylor<3sRin @victoria_lion @chandowd @NaomiBell1997 @dopegyal_

OpenStudy (anonymous):

I don't know what a lemma is.

OpenStudy (anonymous):

Is it lemma or lemon? lol

OpenStudy (usukidoll):

I thought a mod suspended you for mass tagging

OpenStudy (anonymous):

I thought you should get a real education

OpenStudy (uh60blackhawk):

future what grade lvl are you?

OpenStudy (usukidoll):

he's in preschool :)

OpenStudy (anonymous):

Essentially

OpenStudy (abb0t):

You're the worse troll I've ever seen.

OpenStudy (usukidoll):

@abb0t he was suspended for mass tagging...look he's here again

OpenStudy (uh60blackhawk):

what grade lvl are you

OpenStudy (anonymous):

Guys all I want is an answer, please quit spamming

OpenStudy (usukidoll):

oh yay a mod is here ^^ buh bye future whatever

OpenStudy (usukidoll):

We the people can't take FutureMathProfessor seriously due to the massive amounts of tagging on numerous questions.

OpenStudy (abb0t):

Ok, it's not that serious. Lol

OpenStudy (abb0t):

I just banned him, guys. Don't worry. He won't be bothering us for a while.

OpenStudy (uh60blackhawk):

seriously what grade level are?

OpenStudy (usukidoll):

abbot I love you

OpenStudy (compassionate):

Please, stop this.

OpenStudy (anonymous):

you guys are a disgrace lol.

OpenStudy (the_fizicx99):

Lol >.> wut? How'd you ban him?

OpenStudy (compassionate):

@abb0t is an administrator.

OpenStudy (the_fizicx99):

I don't see the purple thingy though ♦.♦

OpenStudy (compassionate):

He's incognito.

OpenStudy (the_fizicx99):

LOL >.>

OpenStudy (uh60blackhawk):

lol

OpenStudy (anonymous):

wow nice thats pretty good at least no more mass taggings lol

OpenStudy (the_fizicx99):

He'll be back, I know he will =.=

OpenStudy (anonymous):

I actually liked being continually tagged rather than controlled by the corrupt adminstration

OpenStudy (the_fizicx99):

^

OpenStudy (solomonzelman):

Wow, @TanteAnna, you really think administration is corrupt? say that to Preetha.

OpenStudy (usukidoll):

ooooooooooo!

OpenStudy (usukidoll):

It's the trolls who ruin the site...

OpenStudy (solomonzelman):

who is the troll that ruins the site?

OpenStudy (the_fizicx99):

Certain people..certain.

OpenStudy (usukidoll):

FutureMathProfessor and the mass tagging that's an example of a troll

OpenStudy (solomonzelman):

I see....

OpenStudy (usukidoll):

It's because of them that I can't really recommend this site too often as I used to.

OpenStudy (the_fizicx99):

^

OpenStudy (usukidoll):

I need chocolate to wash the hot dog taste brb for some candy melts.

OpenStudy (the_fizicx99):

♥_♥

OpenStudy (usukidoll):

I just eat the candy melts ^^

OpenStudy (the_fizicx99):

:O

OpenStudy (solomonzelman):

A drink too?

OpenStudy (the_fizicx99):

Nawww :O *mind blown*

OpenStudy (usukidoll):

prefer pepsi

OpenStudy (anonymous):

u would add like varibles

OpenStudy (anonymous):

if 5x + 5 is even, and since 5 is odd, 5x is also odd 5x = 2s + 1, for some integer s 9x + 3 = 3(3x+1) = 3(3x + 5x-2s) = 3(8x - 2s) = 2 * [3(4x-s)] so it's even

OpenStudy (usukidoll):

no point... poster got banned

OpenStudy (solomonzelman):

(for a reason of course)